OpenStudy (anonymous):

what are the solutions to equation (2x-3)^2-1

OpenStudy (bahrom7893):

equals what?

OpenStudy (anonymous):

i mean equals 1 , sorry

OpenStudy (anonymous):

foil the 2x-3 first.

OpenStudy (anonymous):

(2x-3)(2x-3) 4x^2-12x+9

OpenStudy (anonymous):

then set it equal to zero

OpenStudy (anonymous):

4x^2-12x+9=1 4x^2-12x+8=0

OpenStudy (anonymous):

double checking- u meant =1, not =-1, right?

OpenStudy (anonymous):

yes 1

OpenStudy (anonymous):

Simplify, it means that (2x-3) = 1, or 2x = 4, or x = 2.

OpenStudy (anonymous):

ok . Then u factor 4x^2-12x+8 (factor a 4 out first- makes life easier.)

OpenStudy (anonymous):

so wouldnt it be 1 and 2?

OpenStudy (anonymous):

4(x^2-3x+2) 4(x-2)(x-1) x=2 or x=1

OpenStudy (anonymous):

ur right, tera

OpenStudy (anonymous):

okay :) thanxx for your help